OpenStudy (anonymous):

the sum of three numbers is 12 the first number is twice the sum of the second and third. the third number is 5 less the first.?

OpenStudy (anonymous):

Let the three numbers be x,y and z. x+y+z=12.......(1) x=2(y+z)..........(2) z=x-5..............(3) solve these three equations and get the values of x , y and z

OpenStudy (anonymous):

thanks nitz.

OpenStudy (anonymous):

u are welcome......

OpenStudy (anonymous):

nitz, but what if the answer choices dont have x, y in them there numbers

OpenStudy (anonymous):

what are the choices.........?

OpenStudy (anonymous):

3,6,3 8,1,3 4,1,1 9,1,4

OpenStudy (anonymous):

its B 8,1,3

OpenStudy (anonymous):

thanks nitz :) you rock

OpenStudy (anonymous):

ANY other question....

OpenStudy (anonymous):

yess how do you evaluate the determing by using diagonals?

OpenStudy (anonymous):

-1,2,2 4,-1,1 3,2,2

OpenStudy (anonymous):

sorry i didnt get ur question...........

OpenStudy (anonymous):

evaluate the determinate by using diagonals

OpenStudy (anonymous):

do you get it ?